OpenStudy (anonymous):

Purvi and her friends tipped their waiter a total of twelve dollars for dinners that cost nine dollars, eight dollars, eleven dollars and twelve dollars each. Use a percent proportion to solve for the percent tip they left the waiter. Show your work.

OpenStudy (nurali):

x/100 = $12/($9+$8+$11+$12) 1200=40*x x=30%

OpenStudy (anonymous):

your the best thanks :) @Nurali

OpenStudy (nurali):

My Pleasure.